OpenStudy (anonymous):

Carbon-14 is a radioactive isotope that decays exponentially at a rate of 0.0124 percent a year. How many years will it take for carbon-14 to decay to 10 percent of its original amount?

OpenStudy (wolf1728):

When is the percentage reduced to 10%? Half Life of Carbon-14 = 5,730 years Time = half-life * log(bgng amt/ending amt) / log2 Time = 5,730 * log (100%/10%) / 0.3010299957 Time = 5.730 years * 1/0.3010299957 Time = 19,034.65 years I don't think radioactive decay occurs by percentage per year. .0124% = a factor of .000124 per year That would mean after 8,065 years you are left with nothing? No. After 5,730 years you are left with half After 11,460 years you are left with .25 of what you started with 17,190 0.125 22,920 0.0625 28,650 0.03125 34,380 0.015625 40,110 0.0078125 45,840 0.00390625 51,570 0.001953125 57,300 0.0009765625 If radioactive decay occurred by .0124% per year, that would mean radioactive decay was an arithmetic function which it isn't. It is a geometric function.

OpenStudy (wolf1728):

Here is a webpage with a halflife calculator: http://www.1728.org/halflife.htm